A family purchased a house for $240,000.00 with a down payment of $62,500.00
the rest of the purchase in being financed with a $187.500.00 mortgage at 11%.The
closing date for the sale is 10-day before the beginning of the month, which is the
contractual starting date for the mortgage. What will the lender charge for a 10 day
advance?
Note: Show Principal, Rate and Time when solving.

Answer by Theo(13342)

the simplest calculation would be a simple interest type loan for the 10 day period.

240,000 minus 62,500 = 177,500, not 187,500.

the mortgage amount needs to be 177,500 and not 187,500, unless there are some additional charges in there that do not appear on the surface.

assuming 11% per year, the simple interest rate calculations would be:

177,500 * .11 * 10/365 = 534.93.
or:
187,500 * .11 * 10/365 = 579.35.

however, if the bank is assuming the interest rate is compounded monthly, then the effective annual interest rate is (1 + 11/1200) ^ 12 minus 1 = .115718836.

the simple interest calculations would then be:

177,500 * .115718836 * 10/365 = 562.74
or:
187,500 * .115718836 * 10/365 = 594.45.

using the effective annual interest rate, the compound interest rate calculation would more then likely be:

177,500 * 1.115718836 ^ (10/365) = 533.29.
or:
187,500 * 1.115718836 ^ (10/365) = 563.34.

take your pick.
any of these calculations might apply.
or:
none of these calculations might apply.

this is because it is not known what the particular bank policy is for giving you the money 10 days earlier than the mortgage loan is due to start.

it may be at the same rate, or it may be at a higher rate.
more then likely, it will be at a higher rate.
short term loan interest rates are usually higher than long term loan interest rates.

the answers above assume the bank is only interested in getting the same interest rate that they would have gotten if the mortgage started right away.

it's just a theoretical calculation, not necessarily connected to reality in any way.

the only realistic answer is to check with the bank and see what they would charge for you to get the money 10 days earlier than when the mortgage starts.